A gas in a 5.0 liter balloon at 18°C was taken outside on a hot day. If the temperature outside was 37°C, what was the new volume of the balloon in liters? Round your answer to the tenths place.
1
Expert's answer
2018-03-01T08:47:30-0500
V1/T1=V2/T2

V1=(T2*V1)/T1=((37+273)*5.0/ (18+273)=5.33 l

Answer:5.33 l